# Mixed Signals: A Diverse Point Cloud Dataset for Heterogeneous LiDAR V2X Collaboration

### v2

@@ 3.2.1 Synchronization and Localization
Localization, i.e., estimating vehicle position relative to a global reference frame, is one of the most critical tasks for CAV. To overcome inherent problems of Global Navigation Satellite System (GNSS) in urban environments, we use dense and accurate point cloud maps [31] as references for our localization algorithm. Both the vehicles and the roadside units are localized within a common reference frame, referred to as the $map\_frame$, which serves as the origin of our map. The localization algorithm employs a scan-matching technique [3] to estimate the vehicles’ poses within this map, achieving a maximum positioning error of 15 cm and a heading error of 0.4 degrees. This allows for consistent spatial alignment between the vehicles and the roadside infrastructure. The vehicles’ localization estimates their positions within the $map\_frame$, while the roadside unit is static. We leave details about map construction and usage in Appendix C.2.
@@ 4 Proposed Tasks and Benchmarks
Vehicle AP@ Bike AP@ Pedestrian AP@ Avg. Bandwidth (MB)
@@
Early Fusion 0.71 0.71 0.62 0.62 0.69 0.61 7.79
Attentive Fusion [42] 0.71 0.71 0.61 0.61 0.68 0.59 5.26
@@
F-Cooper [6] 0.78 0.78 0.75 0.62 0.70 0.40 15.31
V2X-ViT [41] 0.85 0.85 0.73 0.66 0.75 0.54 19.36
V2VAM [21] 0.84 0.84 0.81 0.73 0.67 0.48 16.78
where2comm [14] 0.77 0.77 0.71 0.62 0.36 0.16 16.78
@@ 4.2 Benchmark
Evaluation Settings. Since the annotations are made in the coordinate system of TOP, we define the region of interest for the two detection tasks as the range $\left[-51.2,51.2\right]$ meters along both the x and y axes of this coordinate system. For evaluation, we transform objects detected by each agent into this coordinate system. The visibility threshold is set to 5 points. Since timestamp mismatches and localization errors are inherent in real-world applications and consequently present in our dataset, we do not artificially introduce them into the messages exchanged among connected agents (something that is often done in synthetic datasets [42, 41]). We measure object detection performance using Average Precision (AP). Detected objects are matched with ground truth based on their Intersection over Union (IoU) in the bird’s-eye view plane. A detection and a ground truth object are considered a match if their IoU exceeds thresholds of 0.3, 0.5, or 0.7. In addition to AP, we measure the bandwidth consumption of each collaborative method to gauge their practicality. The total bandwidth consumption is calculated by multiplying the number of agents in the collaboration network by the size of the message each agent sends. While the number of agents is not dependent on the collaboration method of choice, the message size is. Therefore, we compute the bandwidth consumption by averaging the size of the messages that agents send, measured in Megabytes (MB). While some intermediate collaboration methods [36, 22, 41] employ specialized compressing algorithms to reduce the message size, other methods [42, 25, 7] do not. To obtain a fair comparison, we measure the size of uncompressed messages.
@@ 4.3.1 Collaborative Object Detection
Domain Gap from prior R.H.T. Datasets. To illustrate the domain gap covered by Mixed Signals, we directly adapt an early-fusion and a late-fusion model trained on the right-hand traffic (R.H.T.) V2X dataset, V2V4Real [43], onto our dataset (grayed-out rows in Table 3). Performance degraded significantly, with vehicle headings often predicted incorrectly, indicating a learned prior from traffic flow (Figure 11). This highlights a substantial domain gap due to left-hand traffic and sensor modality differences, underscoring Mixed Signals’ unique contribution to the V2X perception landscape.
@@ 5 Discussion and Conclusion
Our work presents the Mixed Signals V2X dataset, created through careful data selection, sensor synchronization and localization, and a strong investment in high quality annotations. To the best of our knowledge, our dataset is the first to support heterogeneous sensor configurations with varying positions and orientations, collected in an out-of-domain left-hand traffic country, Australia, providing a diverse dataset addition to the field. We hope that the release of our dataset will facilitate research into complex and realistic settings for V2X perception. Future directions of research include studying communication protocols that ensure both fast transmission and directed communication that targets salient information. More broadly, we hope that our work could potentially further the research in the emerging, yet impactful, field of V2X communication.
@@ Acknowledgement
We thank Runsheng (Derrick) Xu and Hao Xiang for their insightful discussions and moral support throughout the early stage of this project. This research is funded by University of Sydney – Cornell University Ignition Grants/Global Strategic Collaboration Awards, National Science Foundation (IIS-2107161), and the New York Presbyterian Hospital. Minh-Quan Dao is funded by ANNAPOLIS project managed by the French National Agency for Research (ANR-21-CE22-0014), and Katie Luo by AAUW American Dissertation Fellowship.
@@ A.1 Devkit Description
We provide a separate devkit and additionally integrate our dataset into the framework OpenCOOD [42], which offers the implementation of various state-of-the-art collaborative perception methods. As OpenCOOD only provides single-class models, we adapt its implementation of Early, Intermediate, and Late Fusion models to detect three classes, including vehicles, bikes, and pedestrians. We added detection heads of 1-by-1 convolution layers to existing architectures to achieve this. In addition, we add the recent Laly fusion [7] to this framework. Every model in our benchmark uses PointPillar [20] as the backbone. Interested readers can refer to our devkit code releasehttps://github.com/chinitaberrio/Mixed-Signals and extended OpenCOOD integration for further details on architectures and training settings.
@@ B Annotation Instructions
We provide the instructions given to the Segments.aihttps://segments.ai/ annotators in the attached material, titled “Spec Document - Multi-sensor labeling” at the bottom of the appendix. We selected to invest in the quality of the annotations, applying rigorous quality control measures to guarantee accurate and consistent labeled data, minimizing errors, and maintaining high standards.

### v3

@@ 3.2.1 Synchronization and Localization
Localization, i.e., estimating vehicle position relative to a global reference frame, is one of the most critical tasks for CAV. To overcome inherent problems of Global Navigation Satellite System (GNSS) in urban environments, we use dense and accurate point cloud maps [31] as references for our localization algorithm. Both the vehicles and the roadside units are localized within a common reference frame, referred to as the map_framemap\_frame, which serves as the origin of our map. The localization algorithm employs a scan-matching technique [3] to estimate the vehicles’ poses within this map, achieving a maximum positioning error of 15 cm and a heading error of 0.4 degrees. This allows for consistent spatial alignment between the vehicles and the roadside infrastructure. The vehicles’ localization estimates their positions within the map_framemap\_frame, while the roadside unit is static. We leave details about map construction and usage in Appendix C.2.
@@ 4 Proposed Tasks and Benchmarks
Vehicle AP@ Bike AP@ Pedestrian AP@
Avg. Bandwidth
@@
Early Fusion 0.65 0.65 0.65 0.65 0.74 0.67 7.79
Attentive Fusion [42] 0.82 0.82 0.71 0.71 0.74 0.68 5.26
@@
F-Cooper [6] 0.75 0.75 0.68 0.68 0.72 0.65 15.31
V2X-ViT [41] 0.84 0.84 0.71 0.70 0.77 0.70 19.36
V2V-AM [21] 0.83 0.83 0.79 0.79 0.69 0.60 16.78
where2comm [14] 0.77 0.77 0.74 0.74 0.31 0.18 16.78
@@ 4.2 Benchmark
Evaluation Settings. Since the annotations are made in the coordinate system of TOP, we define the region of interest for the two detection tasks as the range $\left[-51.2,51.2\right]$ meters along both the x and y axes of this coordinate system. For evaluation, we transform objects detected by each agent into this coordinate system. The visibility threshold is set to 5 points. Since timestamp mismatches and localization errors are inherent in real-world applications and consequently present in our dataset, we do not artificially introduce them into the messages exchanged among connected agents (something that is often done in synthetic datasets [42, 41]). We measure object detection performance using Average Precision (AP). Detected objects are matched with ground truth based on their Intersection over Union (IoU) in the bird’s-eye view plane. A detection and a ground truth object are considered a match if their IoU exceeds thresholds of 0.3, 0.5, or 0.7. In addition to AP, we measure the bandwidth consumption of each collaborative method to gauge their practicality. The total bandwidth consumption is calculated by multiplying the number of agents in the collaboration network by the size of the message each agent sends. While the number of agents is not dependent on the collaboration method of choice, the message size is. We report the bandwidth consumption by averaging the size of the messages that agents send, measured in Megabytes (MB). While some intermediate collaboration methods [36, 22, 41] employ specialized compressing algorithms to reduce the message size, other methods [42, 25, 7] do not. For fair comparison, we report uncompressed sizes.
@@ 4.3.1 Collaborative Object Detection
Domain Gap from prior R.H.T. Datasets. To illustrate the domain gap covered by Mixed Signals, we directly adapt an early-fusion and a late-fusion model trained on the right-hand traffic (R.H.T.) dataset, V2V4Real [43], onto our dataset (grayed-out rows in Table 3). Performance degraded significantly, with vehicle headings predicted incorrectly, indicating a learned prior from traffic flow (Figure 11). This highlights a substantial domain gap due to left-hand traffic and sensor modality differences, underscoring Mixed Signals’ unique contribution to the V2X perception landscape.
@@ 5 Discussion and Conclusion
Our work presents the Mixed Signals V2X dataset, created through careful data selection, sensor synchronization and localization, and a strong investment in high quality annotations. To the best of our knowledge, our dataset is the first to support heterogeneous sensor configurations with varying positions and orientations, collected in an out-of-domain left-hand traffic country, Australia, providing a diverse dataset addition to the field. We hope that the release of our dataset will facilitate research into complex and realistic settings for V2X perception. Future directions of research include studying communication protocols that ensure both fast transmission and directed communication that targets salient information.
@@ Acknowledgement
We thank Runsheng Xu and Hao Xiang for their insightful discussions and support throughout the early stage of this project. This research is funded by University of Sydney – Cornell University Ignition Grants/Global Strategic Collaboration Awards, National Science Foundation (IIS-2107161), and the New York Presbyterian Hospital. Minh-Quan Dao is funded by ANNAPOLIS project managed by the French National Agency for Research (ANR-21-CE22-0014), and Katie Luo by AAUW American Dissertation Fellowship.
@@ A.1 Devkit Description
We provide a separate devkit and additionally integrate our dataset into the framework OpenCOOD [42], which offers the implementation of various state-of-the-art collaborative perception methods. As OpenCOOD only provides single-class models, we adapt its implementation of Early, Intermediate, and Late Fusion models to detect three classes, including vehicles, bikes, and pedestrians. We added detection heads of 1-by-1 convolution layers to existing architectures to achieve this. In addition, we add the recent Laly fusion [7] to this framework. Every model in our benchmark uses PointPillar [20] as the backbone. Interested readers can refer to our devkithttps://github.com/quan-dao/mixed-signals-devkit and code releasehttps://github.com/acfr/Mixed-Signals-Dataset and extended OpenCOOD integration for further details on architectures and training settings.
@@ B Annotation Instructions
We provide the instructions given to the Segments.ai§§§https://segments.ai/ annotators in the attached material, titled “Spec Document - Multi-sensor labeling” at the bottom of the appendix. We selected to invest in the quality of the annotations, applying rigorous quality control measures to guarantee accurate and consistent labeled data, minimizing errors, and maintaining high standards.